You wish to hit a target from several meters away with a charged coin having a mass of 2.5g ... #1 You wish to hit a target from several meters away with a charged coin having a mass of 2.50 g and a charge of +4500 uC. A downward, uniform electric field of 3.75 N/C exists throughout the region. There is also a uniform magnetic field throughout the region. You aim directly at the target and launch the coin horizontally with a speed of 15.2m/s. The coin travels in a straight line and hits the target (a.) What is the magnitude of the uniform magnetic field? (Include gravity in your calculations.) (b.) What is the direction of the uniform magnetic field?
Answer #1

electrtic frce acting on the coin = Fey = E*q

gravitational force acting Fgy = m*g

net downward force Fy = Fe + Fg

the magnetic force acting Fb = q*v*B

here the charge is moving horizantally with constant speed


the net force = 0

Fb + Fe + m*g = 0

Fb = Fe + mg

q*v*B = (4500*10^-6*3.75)+(2.5*10^-3*9.8)


B = -0.605 T


part(a)

B = 0.605 T


part(b)


from Fb = q*(v ross B)

Fb is in upward direction

so magnetic field is into of the page